### v4

@@ Abstract
Code generation tools are essential to help developers in the software development process. Existing tools often disconnect with the working context, i.e., the code repository, causing the generated code to be not similar to human developers. In this paper, we propose a novel code generation framework, dubbed $A^{3}$-CodGen, to harness information within the code repository to generate code with fewer potential logical errors, code redundancy, and library-induced compatibility issues. We identify three categories of representative information for the code repository: local-aware information from current code file, global-aware information from other code files, and third-party-library information. Results demonstrate that by adopting the $A^{3}$-CodGen framework, we successfully extract, fuse, and feed code repository information into the LLM, generating more accurate, efficient, and highly reusable code. The effectiveness of our framework is further underscored by generating code with a higher reuse rate, compared to human developers. This research contributes significantly to the field of code generation, providing developers with a more powerful tool to address the evolving demands in software development in practice.
@@ I Introduction
Unfortunately, none of existing code generation methods take above three information into consideration simultaneously. As the most common code generation tool, ChatGPTĀ [8] demonstrates commendable coding abilityĀ [17] but relies solely on user input. GitHub CopilotĀ [9], which explicitly utilizes local information spanning the current working context, often struggles to harness the full potential of global information presented in the entire code repository [^1] https://githubnext.com/projects/copilot-view/. When these methods disregard local and global module information, they risk omitting crucial logic during code generationĀ [18]. For instance, they might forget to extract text from a tag before attempting to identify it as a bullet point (is_bulleted_text). Furthermore, neglecting third-party library information may cause the LLMs to randomly select third-party libraries and generate code based on them. If the selected libraries third-party are unsupported under the given environment, it can lead to compatibility issues, such as ā€œModuleNotFoundError: No module named xxx.ā€
@@
Our experiments are driven by two primary objectives: 1) To investigate the effectiveness of three distinct types of code repository information in assisting the model with code generation for specific code repositories. 2) To assess whether our framework can efficiently extract and utilize these three types of information. We crawl and filter 29 representative Python code repositories from PyPI, assembling 13,784 functions. To ensure a diverse benchmark dataset, we conduct a sampling process, selecting representative functions from each repository to total 383 functions. We subsequently label these benchmarks and the 3,830 model outputs generated under different configurations to evaluate the performance of $A^{3}$-CodGen. Our experimental results confirm the efficacy of incorporating three types of code repository-awareness knowledge in aiding the foundational model to reuse code repository knowledge. The $A^{3}$-CodeGen framework excels in recognizing and utilizing these three types of code repository knowledge, leading to exceptional performance metrics in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding code reuse capabilities, our framework is capable of generating concise code that seamlessly integrates with the code repository’s working environment. Furthermore, we identify optimal configurations for local and global information, enabling the model to effectively leverage essential knowledge while avoiding the potential pitfalls of information overload.
@@ II-A Limitation of Existing Code Generation Approaches
    Observation 2: Existing methods focus on generating independent and often small-scale code unitsĀ [35, 16]. According to a previous studyĀ [35], the code generated by LLMs shows lower method-level correctness. Although various benchmarks have been proposed (Details are discussed in the Related Work (SectionĀ VII)), newly generated code still remains questionable.
@@ II-C Bridging the Gap: Making LLMs Aware of Three Types of Repository Knowledge
In the next section, we provide a comprehensive description of how to extract, structure, and utilize local aware, global aware, and third-party library aware information from code repositories.
@@ III-A1 Code Repository Information Extraction
For a code repository, we can effortlessly obtain all code files in the repository’s root folder. Then, for each file, we extract all included functions, gathering both basic code information and semantic code knowledge. This process snowballs to extract all essential repository information.
@@
2) Function Summary. Not every function is elaborated by comments. Even though the comments may exist, their quality varies distinctly. Therefore, we generate the function summary to precisely describe the purpose and functionality.
Function Summarizer Unit. To generate natural language semantic information for a function, we design a prompt as the core of the Function Summarizer unit. This prompt creates summaries for each function. Our Function Summarizer prompt consists of a task description, five demonstration examples, and the input template. The input to this unit is the function’s code, and the output is the model-generated summarization for that function.
@@ III-A2 Third-Party Library Extraction
Third-party libraries often appear in code files, and when specific libraries are provided to a model, it likely has learned how to use them during trainingĀ [43, 44]. Therefore, supplying just the library name allows the model to automatically understand how to use it for a given task. For example, we can ask the model to use the NumPy library to compute the cosine similarity, given the input vector_a and vector_b. In this case, the model can directly call functions in the NumPy library, such as numpy.dot and numpy.linalg.norm, to perform the required computations. The pre-trained model’s prior knowledge and extensive large-scale training enable them to effectively utilize third-party libraries, improving the efficiency and accuracy of code generation. Therefore, we extract the import statements of third-party libraries directly from all code files via AST and saved their library names to construct the third-party library base.
@@ III-B2 Global-Module Knowledge Mining
Figure 3: $A^{3}$ Code Generator Unit
  • [leftmargin=*]
    What-If Code Generator Unit: To retrieve global functions from the function base, we consider textual and code semantic similarity. For code semantic similarity matching, we utilize the code generation capability of the LLM, inputting user requirements to generate a what-if code, which encapsulates the logic to fulfill those requirements and serves as input for code similarity matching in the Global Function Retrieval Unit. For in-context learning, we provide three representative examples. Notably, the generated code is only used for retrieving the global function, so completeness is not required, as it is not the final code.
@@ III-B3 Third-Party Library Knowledge Mining
In the process of code generation, to behavior similar to human developers, it’s crucial for the LLM to be aware of the third-party libraries pre-installed in the current code repository environment. Without this awareness, there’s a risk of generating code that relies on third-party libraries unsupported by the current environment, rendering the code non-functional. To ensure that, we create a comprehensive list of all the third-party libraries found in the third-party library base, as extracted in sectionĀ III-A2. It’s important to note that we extract only the names of these libraries and do not include any of their code. Third-party libraries are publicly accessible and widely applied, so according to our empirical observation, we found that the model has already encountered and learned most of the libraries during its training process. Thus, we only extract the names of these libraries without any of their code. By constructing this list of third-party libraries, we empower the LLM to generate code that aligns with the available environment and utilizes the appropriate libraries effectively, reducing the risk of non-executable code.
@@ III-C $A^{3}$ode Generation
The $A^{3}$ Code Generation Unit, as depicted in Fig.Ā 3, consists of four key components: task description, examples, input, and output. In the task description (purple part), we employ a structured prompt to delineate the repository-aware code generation task. Key components of the prompt include: 1) @persona: Guiding the model to assume the role of an expert developer, providing clarity on the task. 2) @terminology: Defining crucial information in a keyword format. 3) @instruction: This section furnishes explicit instructions for executing the task:
  • [leftmargin=*]
@@
In the @rule, we establish three rules to govern the code generated by the model, including calling import FQN, restricting the use of specified third-party libraries, and ensuring that the generated code is not empty (as the model may occasionally generate abstract functions). We then provide three representative examples. Finally, we input the developer requirement and our 3 knowledge into Unit to get the $A^{3}$ function code (i.e., _to_be_generated_code in Fig.Ā 3 (b)). Notably, developers only needs to input requirements, and all 3 types of repo-aware knowledge are obtained automatically.
@@ IV Experiment Setting
In this section, we outline the default foundation models, dataset, and evaluation metrics for assessing our framework. We introduce four Research Questions (RQs):
  • [leftmargin=*]
@@
    RQ4: What is the performance of our $A^{3}$-CodGen framework?
@@ IV-A Foundation Models
We select GPT-3.5-Turbo-16kĀ [48] as the default foundation LLM for this study due to its excellent performance in various SE tasks, along with its popularity and cost-efficiency. As for the embedding model, we use the ada-embedding-002 modelĀ [49], which has the capability of generating high-quality sentence embeddings, enabling us to perform effective similarity computation and retrieval operations. Similar to other LLM-empowered frameworks in SE domain, the more powerful model is selected, the better performance can be achieved, generally.
@@ IV-B Dataset Collection
While there exist some benchmark datasets designed for code generation, such asĀ [50, 51, 35, 52, 16], they suffer from the following issues: 1) some of the existing datasets contain code repositories that are already present in the training corpora of popular LLMs, leading to an overestimation of the code repository-aware code generation capabilities. 2) existing benchmarks only evaluate class-level or small unit test cases, which is not suitable for our research goals. 3) Based on our literature review, none of existing code generation benchmark datasets can evaluate whether $a^{3}$ CodGen can effectively obtain, understand, and correctly reuse three distinct types of code repository information. Therefore, to address the potential risk of data leakage and accurately reflect the capability of $A^{3}$- CodGen, we created a new benchmark dataset, dubbed RepoEval. This decision stems from our intention to verify that the model truly grasp $A^{3}$ code generation after receiving information we provide.
@@ IV-B1 Code Repository Collection
According to OpenAI’s official website, the training data for GPT-3.5-Turbo-16k is cut off as of September 2021. To ensure that our benchmark dataset contains code repositories that have not been seen during model training, we select projects with the earliest release date after October 1, 2021 on PyPI. We further filter the GitHub repositories by choosing those with more than 100 stars and 100 branches, as these metrics indicate repository quality. Ultimately, we identify 29 code repositories to construct our benchmark, each averaging 82.517 code files.
@@ IV-C Metrics
    Reuse Correctness: After identifying local/global/third-party library information, we assess the correctness of reuse. The metric is determined by the number of exact matches between the functions/libraries reused by the model and those in the benchmark, divided by the total number of functions/libraries reused by the model. For instance, in local reuse, if the model reuses functions $local_{a}$, $local_{b}$, and $local_{c}$, and the benchmark reuses functions $local_{b}$, $local_{c}$, and $local_{d}$, there are 2 exact matches, resulting in a correctness score of 2/3 = 0.667.
@@
Subsequently, we enlist two developers with more than five years of experience in Python programming to conduct independent evaluations. For each generated result, they assess whether the function utilizes local functions, assigning a score of 1 if it does and 0 if it doesn’t. This score contributes to the calculation of reuse awareness. In instances of local function reuse, they identify all the reused functions, forming the basis for calculating reuse correctness. The same process applies to global functions and third-party libraries. After completing their evaluations, they discuss any discrepancies and reach a consensus. In all upcoming experiments, we employ this benchmark for evaluation and annotate the model-generated code using the same method. The Cohen’s Kappa coefficientĀ [56] for annotated data across research questions ranges from 86% to 95%, showing high agreement among assessments.
@@ V Evaluation
V-A The Effectiveness of Local-Aware Knowledge (RQ1)
@@
We use vanilla ChatGPT as a baseline, which does not have any local information in its input. We define four variants of local-aware methods, as shown in TableĀ I.
  • [leftmargin=*]
@@
TABLE I: Comparison of the Local Reuse Awareness results obtained with different Local Information. And find the best configuration for Local Information Extraction.
@@
Vallina ChatGPT N/A N/A N/A 0.715
ChatGPT +Local Functions 0.667 0.514 0.580 0.789
ChatGPT +Local Functions +Class Instance Attributes 0.742 0.633 0.683 0.833
ChatGPT +Local Functions +Class Instance Attributes +Module Variables 0.704 0.523 0.600 0.802
ChatGPT +Local Functions +Class Instance Attributes +Module Variables +Local Module FQN 0.747 0.541 0.628 0.817
V-A1 Result Analysis
TableĀ I shows the reuse awareness results for code generation using vanilla ChatGPT and four types of Local-Aware Knowledge Configurations. Vanilla ChatGPT exhibits a precision and recall of 0 as it does not recognize any local functions. Consequently, in its generated code, new functions are often defined to complete certain logic, resulting in code redundancy.
When given knowledge of Local Functions, the model gains the ability to utilize them, resulting in a Precision of 0.667, Recall of 0.514, F1 score of 0.580, and accuracy of 0.789. This is because knowledge of Local Functions enables the model to directly invoke local functions. Upon additionally providing knowledge of Class Instance Attributes, the model’s Precision, Recall, and F1 score improve by 11.24%, 23.15%, and 17.76%, respectively, compared to providing only knowledge of Local Functions. This improvement is attributed to the provision of Class Instance Attributes, which enables the model to better understand how local functions interact with class variables in the project, leading to more accurate application of local functions.
However, when we introduce the Module Variables, the model’s performance drastically declines, particularly with a 17.38% drop in Recall. This may be due to the increased information load resulting from knowledge of Module Variables, which impacts performance. Nevertheless, the model’s performance is restored and approaches the capabilities of providing knowledge of Local Functions and Class Instance Attributes when we subsequently add knowledge of Local Module FQN. This suggests the existence of subtle interactions within the knowledge configuration. When the model receives only one or two key knowledge types, its performance may be superior to when it receives more information. However, as the volume of knowledge increases to a certain extent, the model’s capabilities strengthen, resulting in improved performance.

### v5

@@ Abstract
LLM-based code generation tools are essential to help developers in the software development process. Existing tools often disconnect with the working context, i.e., the code repository, causing the generated code to be not similar to human developers. In this paper, we propose a novel code generation framework, dubbed $A^{3}$-CodGen, to harness information within the code repository to generate code with fewer potential logical errors, code redundancy, and library-induced compatibility issues. We identify three types of representative information for the code repository: local-aware information from the current code file, global-aware information from other code files, and third-party-library information. Results demonstrate that by adopting the $A^{3}$-CodGen framework, we successfully extract, fuse, and feed code repository information into the LLM, generating more accurate, efficient, and highly reusable code. The effectiveness of our framework is further underscored by generating code with a higher reuse rate, compared to human developers. This research contributes significantly to the field of code generation, providing developers with a more powerful tool to address the evolving demands in software development in practice.
@@ I Introduction
Unfortunately, none of the existing code generation methods take the above three types of information into consideration simultaneously. As the most common code generation tool, ChatGPTĀ [8] demonstrates commendable coding abilityĀ [17] but relies solely on user input. GitHub CopilotĀ [9], which explicitly utilizes local information spanning the current working context, often struggles to harness the full potential of global information presented in the entire code repository. [^1] https://githubnext.com/projects/copilot-view/ When these methods disregard local and global module information, they risk omitting crucial logic during code generationĀ [18]. For instance, they might forget to extract text from a tag before attempting to identify it as a bullet point (is_bulleted_text). Furthermore, neglecting third-party library information may cause the LLMs to randomly select third-party libraries and generate code based on them. If the selected libraries third-party are unsupported under the given environment, it can lead to compatibility issues, such as ā€œModuleNotFoundError: No module named xxx.ā€
@@
Our experiments are driven by two primary objectives: 1) To investigate the effectiveness of three distinct types of code repository information in assisting the model with code generation for specific code repositories. 2) To assess whether our framework can efficiently extract and utilize these three types of information. We crawl and filter 29 representative Python code repositories from PyPI, assembling 13,784 functions. To ensure a diverse benchmark dataset, we conduct a sampling process, selecting representative functions from each repository to total 383 functions. We subsequently label these benchmarks and the model outputs generated under different configurations to evaluate the performance of $A^{3}$-CodGen.
Our experimental results confirm the efficacy of incorporating three types of code repository-awareness knowledge in aiding the foundational model to reuse code repository knowledge. The $A^{3}$-CodeGen framework excels in recognizing and utilizing these three types of code repository knowledge, leading to exceptional performance metrics in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding code reuse capabilities, our framework is capable of generating concise code that seamlessly integrates with the code repository’s working environment. Furthermore, we identify optimal configurations for local and global information, enabling the model to effectively leverage essential knowledge while avoiding the potential pitfalls of information overload. By comparing $A^{3}$-CodGen with GitHub CoPilot, we find that while both perform similarly in local reuse, $A^{3}$-CodGen surpasses CoPilot in global reuse and third-party library reuse. We propose potential methods for improving industry products like CoPilot by incorporating our approach.
@@ II-A Limitation of Existing Code Generation Approaches
    Observation 2: Existing methods focus on generating independent and often small-scale code unitsĀ [34, 16]. According to a previous studyĀ [34], the code generated by LLMs shows lower method-level correctness. Although various benchmarks have been proposed (Details are discussed in the Related Work (SectionĀ VII)), newly generated code remains questionable.
@@ II-C Bridging the Gap: Making LLMs Aware of Three Types of Repository Knowledge
In the next section, we comprehensively describe how to extract, structure, and utilize local aware, global aware, and third-party library aware information from code repositories.
@@ III-A1 Code Repository Information Extraction
For a code repository, we can effortlessly obtain all code files in the entire repository by traversing all the directories within the project. Then, for each file, we extract all included functions, gathering both basic code information and semantic code knowledge. This process snowballs to extract all essential repository information.
@@
2) Function Summary. Not every function is elaborated by comments. Even comments exist, their quality varies; some lack clear descriptions of the function’s basic functionality, while others might be over-detailed. Overly lengthy descriptions can distract from understanding the function’s actual usage and behavior. Therefore, we generate a function summary to precisely describe the purpose and functionality of the function. Previous work has demonstrated that large language models (LLMs) are capable of in-context learningĀ [42, 43, 44], where a task description and examples can enable LLMs to perform well. Thus, we use GPT-3.5-TurboĀ [45] as the foundation model due to its demonstrated code understanding and summarization capabilities, as well as its cost-effectiveness. The selection of the foundation model is flexible, and our method is also applicable to advanced models such as GPT-4Ā [32], although they come at a higher cost.
Figure 3: Function Summarizer Unit
Function Summarizer Unit. To generate natural language semantic information for functions, we design a prompt as the core of the Function Summarizer Unit. This prompt creates summaries for each function. As shown in Fig.Ā 3, the prompt includes a task description (highlighted in purple), five demonstration examples, and an input template. The input to this unit is the function’s code (highlighted in green), and the output is the model-generated summary (highlighted in red).
@@ III-A2 Third-Party Library Extraction
Third-party libraries often appear in code files, and when specific libraries are provided to a model, it likely has learned how to use them during trainingĀ [46, 47]. Therefore, supplying just the library name allows the model to automatically understand how to use it for a given task. For example, we can ask the model to use the NumPy library to compute the cosine similarity, given the input vector_a and vector_b. In this case, the model can directly call functions in the NumPy library, such as numpy.dot and numpy.linalg.norm, to perform the required computations. The pre-trained model’s prior knowledge and extensive large-scale training enable them to effectively utilize third-party libraries, improving the efficiency and accuracy of code generation. Therefore, we extract the import statements of third-party libraries directly from all code files via AST and saved their library names to construct the third-party library base. It is important to note that the same library can have different versions, and API call methods may vary between versions. To address this, the current repository’s Python interpreter or the ā€œpip listā€ command can be used to determine the versions of pre-installed libraries, guiding the model in generating accurate code.
@@ III-B2 Global-Module Knowledge Mining
    What-If Code Generator Unit: In SectionĀ III-A1, we construct a function base derived from the code repository. To retrieve global functions from this base, we consider both textual and code semantic similarity. For code semantic similarity matching, we utilize the code generation capability of the LLM, inputting user requirements to generate a what-if code, which encapsulates the logic to fulfill those requirements and serves as input for code similarity matching in the Global Function Retrieval Unit. For in-context learning, we provide three representative examples. Notably, as the generated code is not the final code and only used to retrieve the global function, so completeness is not required. As shown in the experiment results in SectionĀ V-B, using function descriptions for text similarity and what-if code for code similarity retrieval can retrieve more useful global functions compared to using only function descriptions.
@@ III-B3 Third-Party Library Knowledge Mining
In the process of code generation, to behavior similar to human developers, it’s crucial for the LLM to be aware of the third-party libraries pre-installed in the current code repository environment. Without this awareness, there’s a risk of generating code that relies on third-party libraries unsupported by the current environment, rendering the code non-functional. To ensure that, we create a comprehensive list of all the third-party libraries found in the third-party library base, as extracted in sectionĀ III-A2. It’s important to note that we extract only the names of these libraries and do not include any of their code. This approach is supported by recent studies showing that LLMs can serve as neural knowledge bases for SE knowledge due to their pre-training on extensive source code corpora, which includes third-party library knowledgeĀ [48, 49, 50, 42]. Several works demonstrate library-oriented code generation, where LLMs generate code based on specific library APIsĀ [48, 49, 50, 51]. By constructing this list of third-party libraries, we empower the LLM to generate code that aligns with the available environment and utilizes the appropriate libraries effectively, reducing the risk of non-executable code.
@@ III-C $A^{3}$ode Generation
Figure 4: $A^{3}$ Code Generator Unit
The $A^{3}$ Code Generation Unit, as depicted in Fig.Ā 4, consists of four key components: task description, examples, input, and output. In the task description (purple part), we employ a structured prompt to delineate the repository-aware code generation task. Key components of the prompt include: 1) @persona: Guiding the model to assume the role of an expert developer, providing clarity on the task. 2) @terminology: Defining crucial information in a keyword format. 3) @instruction: This section furnishes explicit instructions for executing the task:
@@
In the @rule, we establish three rules to govern the code generated by the model, including calling import FQN, restricting the use of specified third-party libraries, and ensuring that the generated code is not empty (as the model may occasionally generate abstract functions). We then provide three representative examples. Finally, we input the developer requirement and our 3 knowledge into Unit to get the $A^{3}$ function code (i.e., _to_be_generated_code in Fig.Ā 4 (b)). Notably, developers only needs to input requirements, and all 3 types of repo-aware knowledge are obtained automatically.
@@ IV Experiment Setting
In this section, we first outline the selection of foundation models, the dataset, and evaluation metrics for assessing our framework. We then introduce five Research Questions (RQs):
@@
    RQ4: What is the overall performance of our $A^{3}$-CodGen framework?
    RQ5: How does $A^{3}$-CodGen framework compare to existing code generation tools like GitHub Copilot?
@@ IV-A Foundation Models
We select GPT-3.5-Turbo-16kĀ [45] as the default foundation LLM for this study due to its excellent performance in various SE tasks, along with its popularity and cost-efficiency. As for the embedding model, we use the ada-embedding-002 modelĀ [55], which generates high quality sentence embeddings for effective similarity computation and retrieval operations.Ā [56, 57]. Similar to other LLM-empowered frameworks in SE domain, the more powerful model is selected, the better performance can be achieved, generally.
@@ IV-B Dataset Collection
While there exist some benchmark datasets designed for code generation, such asĀ [58, 59, 34, 60, 16], they suffer from the following issues: 1) some of the existing datasets contain code repositories that are already present in the training corpora of popular LLMs, leading to an overestimation of the code repository-aware code generation capabilitiesĀ [61, 62]. 2) existing benchmarks only evaluate class-level or small unit test cases, which is not suitable for our research goals. 3) Based on our literature review, none of existing code generation benchmark datasets can evaluate whether $A^{3}$-CodGen can effectively obtain, understand, and correctly reuse three distinct types of code repository information. Therefore, to address the potential risk of data leakage and accurately reflect the capability of $A^{3}$-CodGen, we created a new benchmark dataset, dubbed RepoEval. This decision stems from our intention to verify that the model truly grasp $A^{3}$ code generation after receiving information we provide.
@@ IV-B1 Code Repository Collection
To construct a benchmark dataset, we follow methodologies similar to those in previous worksĀ [63, 64, 65]. As we focus on Python code repositories, it is common to collect packages from the PyPI ecosystem, as mentioned in Alfadel et al.Ā [65].
To ensure that the chosen repositories are not part of the training data for the models used, we verify the release dates of each repository to identify their first release date. If the first release date is after the training period of the model, it means that the repository was not included in the training data. According to OpenAI’s official website, the training data for GPT-3.5-Turbo-16k is cut off as of September 2021. Therefore, we select projects with the earliest release date after October 1, 2021, from PyPI. We further filter the GitHub repositories by choosing those with more than 100 stars and 100 branches, as these metrics indicate repository qualityĀ [64]. Ultimately, we identify 29 code repositories to construct our benchmark, each averaging 82.517 code files.
@@ IV-C Metrics
    Reuse Correctness: After identifying local/global/third-party library information, we assess the correctness of reuse. The metric is determined by the number of exact matches between the functions/libraries reused by the model and those in the benchmark, divided by the total number of functions/libraries reused by the model. For instance, in local reuse, if the model reuses functions $local_{a}$, $local_{b}$, and $local_{c}$, and the benchmark reuses functions $local_{b}$, $local_{c}$, and $local_{d}$, there are 2 exact matches, resulting in a correctness score of 2/3 = 0.667. It’s important to note that we do not use conventional metrics like BLEUĀ [69], ROUGEĀ [70], or METEORĀ [71], as they only evaluate token-level similarity between the model-generated code and the benchmark code. These metrics do not effectively measure the model’s ability to correctly reuse code components during generation.
@@
To validate the call relationships between the extracted functions, we recruit two developers with over five years of Python programming experience. Each function in the benchmark dataset is independently labeled by both developers. They evaluate whether the function utilizes local functions within the current code file, assigning a score of 1 if it does and 0 if it doesn’t. Similarly, they assess whether the function utilizes global functions and third-party libraries, assigning scores of 1 or 0 accordingly. These scores contribute to calculating reuse awareness. For functions involving reuse, they identify all reused functions/libraries to calculate reuse correctness.
After the independent labeling, discrepancies are discussed to reach a consensus. If any disagreements remain, the first author resolves them and makes the final decision. In subsequent experiments, we use this benchmark for evaluation and annotate the model-generated code using the same method. To label the model-generated results, we recruit 4 master’s students and 8 undergraduate students, each with at least 2 years of Python programming experience. After a 1-hour training session on our standardized method, tasks are evenly distributed among them. We utilize Cohen’s kappaĀ [72] to assess inter-rater agreement, with coefficients ranging from 86% to 95% for all annotated data (including the benchmark dataset and model-generated results), indicating high agreement among assessments.
@@ V Evaluation
V-A RQ1: Does Local Module Aware Knowledge indeed help LLMs fully utilize local module knowledge during code generation, and which types are most effective in improving local awareness?
V-A1 Motivation
@@
V-A2 Experiment Methodology
We use vanilla GPT-3.5-Turbo-16k as a baseline, which does not have any local information in its input. We define four variants of local-aware methods, as shown in TableĀ I.
@@ V-A2 Experiment Methodology
TABLE I: Comparison of the Local Reuse Awareness results obtained with different Local Information.
@@
Vallina GPT-3.5-Turbo-16k N/A N/A N/A 0.721
GPT-3.5-Turbo-16k +Local Functions 0.667 0.528 0.589 0.796
GPT-3.5-Turbo-16k +Local Functions +Class Instance Attributes 0.720 0.632 0.673 0.830
GPT-3.5-Turbo-16k +Local Functions +Class Instance Attributes +Module Variables 0.728 0.557 0.631 0.820
GPT-3.5-Turbo-16k +Local Functions +Class Instance Attributes +Module Variables +Local Module FQN 0.759 0.566 0.649 0.830
V-A3 Result Analysis
TableĀ I shows the reuse awareness results for code generation using vanilla GPT-3.5-Turbo-16k and four types of Local-Aware Knowledge Configurations. Vanilla GPT-3.5-Turbo-16k exhibits a precision and recall of 0 as it does not recognize any local functions. Consequently, in its generated code, new functions are often defined to complete certain logic, resulting in code redundancy.
When given knowledge of Local Functions, the model achieves a precision of 0.667, recall of 0.528, F1 score of 0.589, and accuracy of 0.796. This is because Local Function knowledge allows the model to directly invoke these functions. Upon additionally providing knowledge of Class Instance Attributes, the model’s precision, recall, and F1 score improve by 7.94%, 19.70%, and 14.26%, respectively, compared to providing only knowledge of Local Functions. This improvement is attributed to the provision of Class Instance Attributes, which enables the model to better understand how local functions interact with class variables in the project, leading to more accurate application of local functions.
However, when we introduce the Module Variables, the model’s performance drastically declines, particularly with a 11.87% drop in precision. This may be due to the increased information load resulting from knowledge of Module Variables, which impacts performance. Nevertheless, the model’s performance is restored and approaches the capabilities of providing knowledge of Local Functions and Class Instance Attributes when we subsequently add knowledge of Local Module FQN. This suggests the existence of subtle interactions within the knowledge configuration. When the model receives only one or two key knowledge types, its performance may be superior to when it receives more information. However, as the volume of knowledge increases to a certain extent, the model’s capabilities strengthen, resulting in improved performance.
We ultimately choose the Local-Aware Knowledge setup, including knowledge of Local Functions and Class Instance Attributes. This configuration serves as default for the upcoming experiments.